OpenStudy (anonymous):

What are the solutions of 2x2 + 16x + 34 = 0? Answer –8 ± 2i –8 ± i –4 ± 2i –4 ± i

OpenStudy (anonymous):

Help please.

OpenStudy (anonymous):

Divide each side by 2.\[x^2+8 x+17=0 \]\[x\text{ = }\frac{-8\pm \sqrt{8^2-4*1*17}}{2*1}=\frac{1}{2} (-8\pm 2 i)=-4\pm i \]